which binomial is a factor of
14aSQUARED-15a+4
A.(-6,-7)
B.(6,7)
C.(-6,-7)
D.(6,-7)

Answer by checkley75(3666)

14A^2-15A+4=0
(7A-4)(2A-1)=0
7A-4=0
7A=4
A=4/7
A=.57 ANSWER.
2A-1=0
2A=1
A=1/2 0R .5 ANSWER.
NONE OF THESE CHOICES ARE CORRECT!!!